You Save Bank has a unique account. If you deposit $5,750 today, the bank will pay you an annual interest rate of 6 percent for 3 years, 6.6 percent for 2 years, and 7.3 percent for 6 years. How much will you have in your account in 11 years?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=5,750*(1.06)^3*(1.066)^2*(1.073)^6

=5,750*2.06552443

=$11876.77(Approx).